A group can charter a particular aircraft at a fixed total cost. If 36 people
charter the aircraft rather than 40 people, then the cost per person is greater by
$12. What is the cost per person if 40 people charter the aircraft?

Answer by ptaylor(2198)

Let x=fixed total cost to charter the plane
Then cost per person for 36 people is x/36
And cost per person for 40 people is x/40
Now we are told that (x/36)-12=(x/40) multiply each term by 360
10x-4320=9x
x=$4320--------------------- total cost to charter the plane
x/40 =4320/40=$108 ---------------cost per person for forty people
4320/36 -12=108
120-12=108
108=108
Hope this helps---ptaylor